OpenStudy (anonymous):

The perimeter of a rectangle is 40 feet. The ratio of the width to the length is 2:3. Find the length and the width

OpenStudy (anonymous):

5x=20 so x=4 lenght: 12 width: 8

OpenStudy (anonymous):

Let length = 2x + 2 Let width = x Let P = 94 P = 2L + 2W 94 = 2(2x + 2) + 2x 94 = 4x + 4 + 2x 94 = 6x + 4 94 - 4 = 6x 90 = 6x 90/6 = x 15 = x The width is 15 feet. The length is 2(15) + 2 or 32 feet. The Proof: 94 = 2(32) + 2(15) 94 = 64 + 30 94 = 94

OpenStudy (anonymous):

@smilekbug your wrong

OpenStudy (anonymous):

getting me back? ;) @ynooran

OpenStudy (anonymous):

where did I go wrong :(

OpenStudy (anonymous):

this is how u do it:

OpenStudy (anonymous):

2:3 so 2x+3x=40/2

OpenStudy (anonymous):

5x=20

OpenStudy (anonymous):

x=4

OpenStudy (anonymous):

so lenght is 3x which is 12

OpenStudy (anonymous):

width is 2x which is 8

OpenStudy (anonymous):

I got it now! I learned something too. Thanks @ynooran :)

OpenStudy (anonymous):

anytime